OpenStudy (anonymous):

make a conjecture about the next item in the sequence. 1, 4, 16, 64, 256 answer.......... a. 122 b. -480 c.-488 d. 116

OpenStudy (anonymous):

umm...none of the above.

OpenStudy (anonymous):

1 = 1^2 4 = 2^2 16 = 4^2 64 = 8^2 256 = 16^2 so next in line should be 32^2...